# Mixing Natural and Synthetic Images for Robust Self-Supervised Representations

### v1

@@ 4 Experiments
Table 3: Top-1 classification accuracies (%) of various models trained on ImageNet-1K, evaluated on domain shift datasets.
Model IN-1K IN-V2 IN-Sketch IN-A IN-R ObjectNet Mean
SimCLR 63.34 50.10 14.08 1.64 23.71 14.64 27.92
SimCLR+Syn 57.58 44.70 16.19 1.72 26.12 11.35 26.28
SimCLR+DiffMix 67.90 54.53 22.57 2.22 34.97 19.65 33.64
Table 4: Comparison of transfer learning performance on six diverse datasets for models trained on ImageNet-1K.
Model CIFAR-10 CIFAR-100 Aircraft DTD Flowers102 Food101 Mean
SimCLR 72.80 47.11 18.99 57.28 73.13 64.05 55.56
SimCLR+Syn 71.04 44.36 17.94 51.80 69.02 58.06 52.04
SimCLR+DiffMix 77.90 53.02 24.36 58.59 82.94 66.40 60.53
Model Corrupted Clean Mean
SimCLR 18.87 30.16 24.96
SimCLR+Syn 18.79 28.11 23.25
SimCLR+DiffMix 23.94 35.02 29.93
Training Datasets.
The substantial size of the ImageNet-1K (IN-1K) (Deng et al., 2009) dataset, which contains approximately 1.3 million images, presents challenges for extensive experimentation. Consequently, we primarily utilize the more manageable ImageNet-100 (IN-100) dataset (Tian et al., 2019) for our studies, which include 100 classes and 1300 images per class. This dataset’s smaller scale enables us to efficiently run multiple variations of each synthetic dataset and thoroughly evaluate the impact of various design choices. Nonetheless, we extend our experiments to IN-1K with SimCLR to validate our findings on a larger scale.
4.1 Robustness to distribution shifts
To evaluate performance under domain shifts, we selected a variety of datasets including ImageNet-A (IN-A) (Hendrycks et al., 2019), ImageNet-Sketch (IN-Sketch) (Wang et al., 2019), ObjectNet (Barbu et al., 2019), ImageNet-V2 (IN-V2) (Recht et al., 2019), and ImageNet-R (IN-R) (Hendrycks et al., 2020). We trained variants of DINO, SimCLR, and Barlow Twins models using real, synthetic, and a blend of both image types from the IN-100 dataset.
A key aspect of our methodology is the training of linear probes: they are consistently trained on real images from the same dataset as the model. For instance, models trained with synthetic IN-100 images use real IN-100 images for linear probe training. To train models on IN-100 we used subsets of the datasets with common categories with IN-100 (ObjectNet was excluded due to lack of common labels with IN-100).
Models leveraging DiffMix exhibit higher accuracy on the IN-100 and other datasets, compared to models trained solely on real or synthetic images (Table 1). This indicates that DiffMix enhances robustness to natural distribution shifts.
In Table 3, we drill down on the SimCLR variants on the IN-1K dataset, and these findings align with our previous observations from the IN-100 dataset. While DINO outperforms on IN-100, we chose SimCLR for this experiment owing to its markedly quicker training time, attributed to its use of fewer crops and simpler data augmentations.
Notably, the model trained with DiffMix exhibits superior performance compared to those trained solely on either type. This advantage is consistent across various natural distribution datasets. However, it is crucial to highlight that accuracy on more challenging datasets like ImageNet-A is still low (Tomasev et al., 2022; Djolonga et al., 2020). This may be attributed to the backbone of the SimCLR model being ResNet-50, while ImageNet-A was curated specifically as images that fool the ResNet-50 model (Hendrycks et al., 2019).
4.2 Domain transfer on different datasets
Our models’ feature generality was assessed by testing their performance across different image domains, using a wide range of datasets to evaluate DiffMix’s effectiveness. The datasets include: Aircraft (Maji et al., 2013), DTD (Describable Textures Dataset) (Cimpoi et al., 2013), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), CIFAR-10 (Krizhevsky, 2009), and CIFAR-100 (Krizhevsky, 2009).
We leveraged the SimCLR model, training it on a mix of real, synthetic, and DiffMix data within the IN-1K dataset. For each variant, we first pre-trained SimCLR’s backbone, subsequently freezing these pre-trained layers to focus on training linear probes. The top-1 accuracy results for each dataset variation are presented in Table 4. These findings demonstrate our approach’s superiority in achieving higher top-1 accuracy across all datasets, surpassing models trained solely on real or synthetic images. For a fair comparison, we standardized the evaluation by using the same SimCLR hyperparameters for all variations. Although our training iterations are lower (100 epochs) as compared to the original SimCLR (1000 epochs), resulting in slightly lower numbers, our method still consistently outperforms the original SimCLR model in relative terms.
4.3 Performance on low-quality images
In this section, we assess the performance of our pre-trained models on the IN-1K dataset, using the VizWiz-Classification (VW-C) dataset (Bafghi & Gurari, 2023). This dataset is notable for containing images taken by people who are blind with various quality issues, such as blur, underexposure, overexposure, poor framing, obscured, and rotated views. We evaluate model performance on corrupted (the mean accuracy of six quality issues), clean, and all images of the dataset.
Table 5 outlines our results. Interestingly, the model trained only on synthetic data, which typically have lower quality, exhibits a smaller drop in accuracy on corrupted images (0.08%) than seen with the clean images of VW-C (2.05%) and the ImageNet validation dataset (6.06%). Finally, our framework can enhance SimCLR models’ ability to manage images with quality issues, highlighting its practical value in real-world scenarios where image quality can vary.
Figure 3: Comparison of top-1 accuracy on IN-100 for SimCLR models trained with and without DiffMix across varying scales of training images. The green shaded area indicates the performance gap between the two models.
4.4 Training with limited data
We explored the impact of limited data on SSL, assessing DiffMix’s effectiveness. Using 25%, 50%, and 100% of the IN-100 dataset and synthetic images generated with a guidance scale of 8, we trained SimCLR models on both original and DiffMix-augmented images. Linear probes were trained on the real images of each subset. Results show that performance gains persist, with a 5.14% increase even when training on only 25% of the images.
5 Ablation study of DiffMix
In this section, we empirically study DiffMix applied to SimCLR and DINO, focusing on finding the optimal guidance scale, understanding the effect of mixing levels on accuracy and self-attention mask, and evaluating the impact of data augmentation on model performance.
5.1 Existence of optimal guidance scale
Figure 4: Left: Distribution of mean cosine distance between feature vectors of images of all categories of IN-100 and generated images with different guidance scales. Right: The relationship between guidance scales used in image generation and the corresponding FID scores of the generated images (Blue line). Additionally, it shows the top-1 accuracy of SimCLR models when trained only on synthetic images, evaluated on the ImageNet-100 dataset (Green line).
In our study, we aimed to identify the optimal guidance scale for generating images on IN-100, which could then be applied to larger-scale image generation for IN-1K. This follows the methodology outlined in (Tian et al., 2023). As demonstrated in Figure 4, we found that using a guidance scale of 8 yields the highest top-1 accuracy on the IN-100 validation dataset. This finding is consistent with the results from (Tian et al., 2023), despite our use of different generative models and image generation approaches. We plot the Frechet Inception Distance (FID) scores (Heusel et al., 2017) to assess the similarity between generated images and IN-100 source images across guidance scales. Interestingly, in general, we observe an inverse relationship between accuracy and image quality. Additionally, in the Left Figure 4, following (Bafghi & Gurari, 2023), we calculate the mean cosine distances between all feature vector pairs per category using CLIP (Radford et al., 2021) for feature extraction, demonstrating that generated images offer more diversity than real images. This diversity may be attributed to the randomness introduced by our generative model.
5.2 DINO: Impact of mixing in self-attention
@@
In our analysis of DINO (Caron et al., 2021) models trained on IN-100, we interpret masks derived from self-attention maps by applying thresholds for enhanced visualization. These maps, sourced from the top-performing head of each ViT-S/16 (Dosovitskiy et al., 2020) DINO model trained on both real and synthetic ImageNet datasets using our DiffMix approach, are not designed for mask creation but rather to highlight the model’s focus areas during image processing. Our findings, illustrated in Figure 5, show that models trained on real images effectively segment objects with some background attention. In contrast, the model trained on synthetic images shows a tendency to focus less on the background, but the overall object segmentation appears somewhat less defined. The DiffMix-trained model strikes a balance, demonstrating clearer object focus with minimal background distraction, indicating improved object segmentation. This improvement is clearly visible in the last row of Figure 5. This suggests that integrating synthetic data with the DiffMix method may enhance scene understanding and image segmentation capabilities.
Figure 5: We present the masks generated by applying a 65% threshold to the self-attention maps. The above maps were obtained using the best head from each ViT-S/16 DINO model that has been trained on ImageNet-100 using real and synthetically generated ImageNet versus DiffMix.
5.3 Performance trend with varying mixing levels
@@
We explore various configurations of image splitting between real and synthetic images for the DINO model. DINO utilizes a total of 10 crops: the teacher model processes two global crops, while the student model analyzes eight local views in addition to the global views. Figure 2 presents a comparison of different configurations and their corresponding top-1 accuracy on the IN-100 validation dataset. All synthetic images are generated with a guidance scale of 8.
@@
Figure 6: Top-1 accuracy comparison of SimCLR models trained with and without DiffMix, and the impact of including versus omitting SSL data augmentations.
5.4 Impact of Data Augmentation
Our study investigates the effect of data augmentation on the SimCLR model, comparing its performance when trained with DiffMix versus solely real images of IN-100. Specifically, for models trained without augmentation, we retain only the random flip and eliminate all other augmentations.
As depicted in Figure 6, an interesting finding is that omitting data augmentations—a key component in self-supervised learning models—does not significantly affect the performance of models trained with both real and synthetic images. The DiffMix model without augmentation generally surpasses the original SimCLR, except for the ImageNet-Sketch dataset. Conversely, In stark contrast, SimCLR models trained exclusively on real images experience a pronounced accuracy decrease in the absence of data augmentations. Specifically, the average accuracy drop across five validation datasets is 15.52% for SimCLR without augmentation, compared to a lesser accuracy drop of 6.60% for SimCLR+DiffMix. Notably, removing these augmentations can lead to faster training times, which is significant considering that data augmentations are often a major bottleneck in the training process, as highlighted in (Bordes et al., 2023).
6 Conclusion
This work demonstrates the potential of mixing synthetic images generated without any labels, with real images in self-supervised pretraining. When used across existing SSL techniques, DiffMix consistently improves the methods across various benchmarks, without any extra data. We identify several properties that could be exploited in future works leveraging synthetic images in SSL. Lower-quality synthetic images from Stable Diffusion can lead to better representations when compared with higher-quality images of the same resolution. Mixing synthetic data with real images can serve as an alternative to augmentations in current SSL methods, proving especially beneficial in domains where applying augmentations is challenging. We further discover that mixing of synthetic data in models like DINO could result in less emphasis on the background, hinting at an enhanced understanding of scene layout within the features that can be leveraged for image segmentation.
Our research reveals a noticeable performance gap between models trained solely on synthetic images and those trained on actual images. This indicates a significant unexplored potential in the development of generative models to improve efficiency and produce images that more closely mimic the distribution of real images. In future work we aim to investigate advanced generative diffusion models that could facilitate better mixing and enhance the robustness of visual features.
Appendix A Overview
  1. 1.
    Provides additional insights into the methodology (complements Section 3).
  2. 2.
Appendix B Background: Joint-embedding SSL
One of the most successful self-supervised learning paradigms is joint-embedding SSL. Joint-Embedding Self-Supervised Learning (SSL) operates on the concept of instance-based views, where each image is treated as a distinct class (Dosovitskiy et al., 2014; Assran et al., 2023). This methodology capitalizes on data augmentation to generate diverse views of the same image. A notable subset of this approach is contrastive SSL techniques(Oord et al., 2018; He et al., 2020; Chen et al., 2020), which, within joint-embedding frameworks, aim to closely align the output embeddings of an image with those of its augmented version. Concurrently, these techniques strive to differentiate these embeddings from those of other images and their respective augmentations. Such methods are commonly implemented in Siamese network architectures (Chen & He, 2021), characterized by two parallel branches that are identical in structure and share the same weights. In our research, we focus on two joint-embedding frameworks: SimCLR(Chen et al., 2020) and BarlowTwins(Zbontar et al., 2021).
B.1 Training and evaluating Joint-embedding SSL
Appendix C Method
C.1 Mixing in Distillation SSL
Recent methods such as DINO (Caron et al., 2021), exploit architectural tricks inspired by knowledge distillation (KD) (Hinton et al., 2015). The KD technique involves training a ‘student’ network to emulate the output representations of a ‘teacher’ network. The teacher network’s weights are dynamically updated as a running average of the student network’s weights or are shared with the student network (Grill et al., 2020). Notably, in this configuration, gradients are not backpropagated through the teacher network (Chen & He, 2021).
DINO and DINO + DiffMix:
The original implementation of DINO (Caron et al., 2021) employs a multi-crop strategy to process multiple scaled views of an image. Specifically, for a given image, a set $V$ of different views is generated, comprising two global views ($x^{g}_{1}$ and $x^{g}_{2}$) at higher resolutions and multiple local views ($x^{l}_{k}$ for $k$ ranging from 1 to 8) at lower resolutions. Consider he training framework for DINO leverages the following KD framework: a student network $g_{\theta_{s}}$, parameterized by $\theta_{s}$, is trained to align with the output of a teacher network $g_{\theta_{t}}$, parameterized by $\theta_{t}$. All views (local + global) are processed by the student, while only the global views are handled by the teacher, promoting a ’local-to-global’ learning correspondence.
Building on this foundation, we introduce image mixing in DINO, termed DINO + DiffMix. In this modified approach, instead of using two global views and eight local views from each real image, we employ a mix of real and synthetic images. Specifically, we use one global view and six local views from a real image, supplemented by one global view and two local views from a corresponding synthetic image as shown in Figure 2 (C). This particular mix of real and synthetic views was determined empirically, and we explore varying levels of mixing and its impact on DINO model performance in Section 5.3. Consequently, our modified set includes a global view from a real image ($x^{g}_{1}$), a global view from a synthetic image ($\tilde{x}^{g}_{2}$), and four local views each from the real ($x^{l}_{k}\forall k\in[1,4]$) and synthetic ($\tilde{x}^{l}_{k}\forall k\in[1,4]$) images. In this framework, both the student and teacher networks generate probability distributions over K dimensions, denoted as $P_{s}$ and $P_{t}$, respectively. These probabilities are derived by normalizing the networks’ outputs using a softmax function. Specifically,
Appendix D Experiments
D.1 Training configurations
Our training employs FFCV-SSL (Bordes et al., 2023), an extended version of the FFCV library (Leclerc et al., 2023) with added support for SSL data augmentations, enhancing training speed and feasibility within our limited resources. We have adapted the code for the DINO framework, in addition to Barlow Twins and SimCLR models. For these models, the LARS (You et al., 2017) optimizer with a learning rate of 1 is used, while DINO models utilize the AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of $0.0005$. All experiments are conducted with a consistent batch size of 256, employing the Cosine Decay scheduler (Loshchilov & Hutter, 2016). Offline linear probes are conducted with the AdamW optimizer at a learning rate of 0.0001 across all experiments, utilizing only real images for training in every case.
D.2 Model configurations
In our experiments, SimCLR and Barlow Twins utilize ResNet-50 as the base encoder, while DINO employs Vision Transformer Small/16 (ViT-S/16) as its backbone. Additionally, a linear layer mapping the representation size to the class count is added atop the encoder. The representation size for SimCLR and Barlow Twins is 2048, whereas for DINO, it’s set to 65536, identified as optimal for DINO (Caron et al., 2021).
D.3 Generating images
Our process for generating images utilizes 50 diffusion steps, as noted in (Sariyildiz et al., 2022), expanding the step count beyond 50 appears to be unnecessary. For the IN-100 dataset, we use various guidance scales (${2,3,6,8,12}$) to generate images. Following the findings of (Azizi et al., 2023), we produce large images with a resolution of $512\times 512$ pixels, batched in groups of 64. For the IN-1K dataset, a consistent guidance scale of 8 is applied. We set the random seed to 25 for reproducibility. By generating images before training, we reduce the overall training time.

### v2

@@ D.2 MixDiff boosts robustness to distribution shifts
Appendix E Ablation Study
E.1 DINO: Impact of mixing in self-attention and performance
@@
In our analysis of DINO [12] models trained on IN-100, we interpret masks derived from self-attention maps by applying thresholds for enhanced visualization. These maps, sourced from the top-performing head of each ViT-S/16 [21] DINO model trained on both real and synthetic ImageNet datasets using our MixDiff approach, are not designed for mask creation but rather to highlight the model’s focus areas during image processing. Our findings, illustrated in Figure 10, show that models trained on real images effectively segment objects with some background attention. In contrast, the model trained on synthetic images shows a tendency to focus less on the background, but the overall object segmentation appears somewhat less defined. The MixDiff-trained model strikes a balance, demonstrating clearer object focus with minimal background distraction, indicating improved object segmentation. This improvement is clearly visible in the last row of Figure 10. This suggests that integrating synthetic data with the MixDiff method may enhance scene understanding and image segmentation capabilities.
E.2 Performance trend with varying mixing levels
@@
We explore various configurations of image splitting between real and synthetic images for the DINO model. DINO utilizes a total of 10 crops: the teacher model processes two global crops, while the student model analyzes eight local views in addition to the global views. Table 7 presents a comparison of different configurations and their corresponding top-1 accuracy on the IN-100 validation dataset. All synthetic images are generated with a guidance scale of 8.
@@
E.3 Impact of data augmentation
This experiment examines various MixDiff configurations, including mixing, sequential, and versatile diffusion (definition borrowed from Section 5.1), comparing them against supervised baselines. We use a ResNet-50 architecture for our supervised baseline, training it with standard, synthetic, and mixing approaches. Our evaluation covers both the ImageNet-100 (IN-100) test set and several distribution shift datasets. Our findings reveal that the standard supervised ResNet-50 model performs best on the IN-100 test set but slightly underperforms on distribution shift datasets. Interestingly, the SimCLR model trained with MixDiff (versatile diffusion) achieves the highest average performance across distribution shift datasets, suggesting enhanced generalization capabilities. The supervised model trained with mixing shows the best performance on ImageNet-A (IN-A), possibly due to the dataset’s synthetic characteristics. The SimCLR model with MixDiff (versatile diffusion) also demonstrates comparable performance on IN-A.
E.4 Additional baselines
We take variants of MixDiff like mixing, sequential and versatile diffusion borrowed from section 5.1 and compare them with supervised baselines. For supervised ResNet-50, we use train and evaluate all three variants (standard, synthetic and mixing). Table 5 summarizes these results. We observe that while standard supervised has the best performance on IN-100 test set, it lags a little behind in distribution shift datasets. SimCLR trained with MixDiff (versatile diffusion) outperforms both SimCLR and SimCLR+Syn, demonstrating the effectiveness of our method using different generative model.
E.5 Computational Costs
Figure 11: Total compute time trend comparison with increasing epochs for SimCLR, SimCLR+MixDiff with and w/o augmentations, respectively.
One limitation of using synthetic data is computation time. However, as discussed in Section 4.3, our method is beneficial in limited data scenarios and needs only one generated image per real image. Incorporating MixDiff into SSL pre-training initially slows it down due to synthetic image generation. We compare SimCLR’s training time with augmentations to SimCLR+MixDiff without augmentations on an NVIDIA A100 GPU with a batch size of 256. Using MixDiff without additional augmentations results in only a minimal performance drop, suggesting that synthetic images can complement natural images without needing augmentations. This synthetic image generation process is a one-time upfront cost (amount of time shown in red), unlike traditional augmentations that incur computational costs at each training iteration. Once synthetic images are generated, training MixDiff without any augmentations is super efficient. This gap can be seen as the difference between the circular grey point and the lower diamond green point. Figure 11 shows that MixDiff’s one-time costs scale better compared to SimCLR, whose augmentations incur costs at each iteration. We recommend using MixDiff with traditional augmentations for the best performance. The enhanced quality of learned representations and improved generalization across diverse scenarios justify the computational trade-off, especially where robustness and adaptability are crucial.